OpenStudy (anonymous):

calvin works as an air conditioning contractor. the customer wants to buy an air conditioner for his family room. his room measures 30 ft. by 16 ft. his walls are 8ft. high. to determine how large a unit to buy, he must calculate the volume of the room. how many cubic feet of air are in the room? @smileyxl3

OpenStudy (smileyxl3):

Formula \[V = l \times w \times h\] So \[V = 30 \times 16 \times 8 =\]